OpenStudy (anonymous):

Help i need to simplify 8+√(128)/32 is it 1+√(2)/4 or 1/32(8+8√(2)

OpenStudy (dumbcow):

\[8+\frac{\sqrt{128}}{32}\] is this correct

OpenStudy (anonymous):

yes

OpenStudy (dumbcow):

\[=8 + \frac{8\sqrt{2}}{32}\] \[=8 + \frac{\sqrt{2}}{4}\]

OpenStudy (anonymous):

thx u